a get together party charged at $8/head became $9/head because 8 will not attend the party. how many people attended the party?
note: this question was printed and given as a handout in class so i do not know the book reference
thank you

Answer by stanbon(75887)

a get together party charged at $8/head became $9/head because 8 will not attend the party. how many people attended the party?
-----------------
Original Plan:
Let number in the party be "n":
Let total cost of the party be "C"
---
Then average cost per individula is C/n = 8
-----------------------------
New Plan:
Number in the party is "n-8"
Total cost of the party is still "C".
Then average cost per individual if C/(n-8) = 9
----------------------
Equation:
From original you get C = 8n
Substitute into new plan to solve for "n":
8n/(n-8) = 9
8n = 9n-72
n = 72 (original number in the party)
n-8 = 64 (number that actually attended the party)
==========================================
Cheers,
Stan H.